A glass bottle of interval volume is treated from 20°C - 50°C. If the linear expansivity of the glass is 9×10–4, the volume of the bottle as 50°C is?


1
Expert's answer
2021-06-14T12:19:37-0400
"V'=V_0(1+\alpha\Delta T)^3\\
V'=20000(1+9\cdot10^{-6}(50-20))^3\\=20016.2\ cm^3"